OpenStudy (anonymous):

How does the zero product property help you find the x-intercept of y = x² - x - 12?

OpenStudy (hunus):

y = x^2 - x - 12 = (x - 4)(x + 3) Now you have your x intercept when y is equal to zero The zero product property says that if ab = 0 then either a or b or both equal zero. So here when (x - 4) or (x - 3) equal zero you have your x intercept

OpenStudy (hunus):

Do you still need help?